# Scaling Laws of RoPE-based Extrapolation

### v1

@@ Abstract
The extrapolation capability of Large Language Models (LLMs) based on Rotary Position Embedding Su et al. (2021) is currently a topic of considerable interest. The mainstream approach to addressing extrapolation with LLMs involves modifying RoPE by replacing 10000, the rotary base of $\theta_{n}={10000}^{-2n/d}$ in the original RoPE, with a larger value and providing longer fine-tuning text. In this work, we first observe that fine-tuning a RoPE-based LLM with either a smaller or larger base in pre-training context length could significantly enhance its extrapolation performance. After that, we propose Scaling Laws of RoPE-based Extrapolation, a unified framework from the periodic perspective, to describe the relationship between the extrapolation performance and base value as well as tuning context length. In this process, we also explain the origin of the RoPE-based extrapolation issue by critical dimension for extrapolation. Besides these observations and analyses, we achieve extrapolation up to 1 million context length within only 16K training length on LLaMA2 7B and 13B (Touvron et al., 2023b).
@@ 1 Introduction
Large Language Models (LLMs) have become the dominant architecture in a variety of natural language processing tasks(OpenAI, 2023; Touvron et al., 2023a; b), while Transformers (Vaswani et al., 2017) based on Rotary Position Embedding (RoPE) (Su et al., 2021) have become the dominant backbone in wide range of LLM design (Chowdhery et al., 2022; Nijkamp et al., 2022; Touvron et al., 2023a; b). While RoPE can theoretically represent sequences through trigonometric functions, as detailed in Appendix A, its performance drops when the input sequence or context length surpasses the training length(Press et al., 2021; Chen et al., 2023), seen in Figure 1. This extrapolation problem (Press et al., 2021) limits tasks like long text modeling and summarization(An et al., 2023).
@@
Figure 1: Perplexity on the validation data of Books3 (Presser, 2020) with different extrapolation methods, including Dynamic NTK (LocalLLaMA, 2023b). RoPE fine-tuned with either a smaller or larger base on the original training length of 4K or a much longer context of 16K, could outperform other extrapolation strategies and extrapolate to 100K context length.
Concerning the extrapolation issue with RoPE, different works have provided various interpretations and corresponding solving attempts. These works could divided into two schools of thought. One limits the scope of self-attention (Ratner et al., 2022; Han et al., 2023) given the fact that self-attention computations in RoPE fail to keep stable beyond training context and exhibit attention score explosion as well as monotonous entropy increase (Chen et al., 2023; Han et al., 2023). The other aims to capture longer contexts by using smaller rotation angles and longer fine-tuning context (Chen et al., 2023; Peng et al., 2023). Currently, popular methods, such as Dynamic NTK (LocalLLaMA, 2023a) and Code LLaMA (Rozière et al., 2023), mainly come from the second approach. Both approaches adapt RoPE to longer contexts with a larger rotary base. Specifically, Dynamic NTK (LocalLLaMA, 2023a) adjusts the base with a coefficient increasing with the length of inference, allowing RoPE-based LLMs to adapt simultaneously to longer context, while Code LLaMA (Rozière et al., 2023) directly sets the base at 1000000 and gets further trained on sequences with 16K tokens, yielding a context beyond 100K.
@@
In this paper, we conduct further experiments on increasing and decreasing the rotary base in Section 2 and subsequently discover that adjusting the rotary base in both directions can contribute to the extrapolation of RoPE-based LLMs. Building upon these observations, we provide a comprehensive explanation for the seemingly counter-intuitive phenomenon from a periodic perspective. Meanwhile, we establish a unified theoretical framework for RoPE-based extrapolation known as the Scaling Laws of RoPE-based Extrapolation [^1] It is important to note that the scaling laws proposed in this work are irrelevant with the well-known scaling laws (Kaplan et al., 2020). In this paper, scale refers to the adjustment of RoPE’s rotary angles (LocalLLaMA, 2023a), rather than the change of the model size. Unlike the scaling laws (Kaplan et al., 2020) that are empirically derived, our scaling laws define mathematical relations between context window size and rotary base, supported by experiments.. We pinpoint specific changes during base reduction that lead to a significant boost in extrapolation in Section 3.1 and identify the upper bound of extrapolation for larger bases in Section 3.3. This clarifies how Code LLaMA (Rozière et al., 2023) manages a 100K extrapolation with only 16K training context. Furthermore, we validate our theories in Section 3.4 and Appendix C, shedding light on both the core principles of Dynamic NTK (LocalLLaMA, 2023a) and pinpointing instability sources self-attention computations in RoPE-based extrapolation (Han et al., 2023). Finally, we present the contributions and guiding significance of this work for other methods that achieve extrapolation during the inference phase. In summary, our contribution can be summarized as follows:
@@ 2.1 Larger Bases Promise Better Extrapolation
First, Larger bases allow LLaMA2 (Touvron et al., 2023b) to extrapolate beyond its training context length, aligning with findings from Rozière et al. (2023). Secondly, the extrapolation with larger bases has a clear limit where language modeling perplexity stays consistent. Beyond this limit, the extrapolation performance declines significantly. Furthermore, as the base value rises, LLaMA2 can extrapolate to a longer context. Finally, compared to Dynamic NTK (LocalLLaMA, 2023a), RoPE tuned with larger bases degrades much quicker beyond its extrapolation upper bound. Therefore, for fine-tuning with larger bases, the performance beyond the upper bound could be consistently overtaken by Dynamic NTK. Nevertheless, within the upper bound, this approach still outperforms Dynamic NTK by a considerable margin, leading to a context beyond 100K with only a 4K tuning length, when the base is set over 600000.
2.2 Smaller Bases Also Promise Better Extrapolation
We then conduct the extrapolation experiments with smaller bases, using the same setup as for larger bases. Interestingly, even though this goes against common research findings (LocalLLaMA, 2023a; Rozière et al., 2023), fine-tuning RoPE with smaller bases on the original context length still boosts extrapolation, as shown in Figure 3. It also extends the context window beyond the training length. Yet, there are distinct differences when comparing RoPE with smaller bases to larger ones.
(a) Results on LLaMA2 7B
(b) Results on LLaMA2 13B
Firstly, RoPE with smaller bases does not have a distinct upper bound of extrapolation. While perplexity worsens as context length grows, this decline is gentler with smaller bases. Secondly, the enhancement in extrapolation of RoPE with smaller bases is not uniform. Between a base of 10000 and 8000, the extrapolation performance exhibits a tiny improvement. Then between a base of 8000 and 2608, improvement is moderate. After that, from 2608 to 1304 and further to 652, the improvement becomes more pronounced. Finally, when the rotary base is 500, the extrapolation curve becomes sufficiently smooth thus resulting in strong extrapolation over 48K context length and superior performance over Dynamic NTK (LocalLLaMA, 2023a).
(a) Results on LLaMA2 7B
(b) Results on LLaMA2 13B
Figure 4: Perplexity and accuracy on the validation data from Books3 dataset (Presser, 2020) of LLaMA2 7B and 13B (Touvron et al., 2023b). In (a) and (b), the first row shows perplexity while the second row shows accuracy, both measured cumulatively. The first column shows the change of extrapolation performance w.r.t. rotary base at different context windows with a line plot, while the second column visualizes it with a heat map. In the heat map of perplexity, the value is log-scaled, namely, cross-entropy loss of language modeling.
@@ 3.2 Critical Dimension for Extrapolation
$\bm{A}_{t,s}&=\mathrm{Re}\left[\underbrace{\color[rgb]{.75,0,.25}\sum_{n=0}^{d/2-1}\tilde{q}_{t}^{(n)}\tilde{k}_{s}^{(n)}{}^{*}e^{i(t-s)\theta_{n}}}_{\text{full attention scores in RoPE}}\right]\\ &=\mathrm{Re}\left[{\underbrace{\color[rgb]{0,0,1}\sum_{n=0}^{d_{\text{extra}}/2-1}\tilde{q}_{t}^{(n)}\tilde{k}_{s}^{(n)}{}^{*}e^{i(t-s)\theta_{n}}}_{\text{reliable part for extrapolation}}+\underbrace{\color[rgb]{1,0,0}\sum_{n=d_{\text{extra}}/2}^{d/2-1}\tilde{q}_{t}^{(n)}\tilde{k}_{s}^{(n)}{}^{*}e^{i(t-s)\theta_{n}}}_{\text{OOD part for extrapolation}}}\right]\end{aligned}\text{.}$ (2)
@@
$T_{n}=\frac{2\pi}{\theta_{n}}=2\pi\cdot{10000}^{\frac{2n}{d}}\leq T_{\text{train}}\text{,}$  for $\text{\quad for\ }n=0,\cdots,d_{\text{extra}}/2-1\text{,}$ (3)
$T_{n}=\frac{2\pi}{\theta_{n}}=2\pi\cdot{10000}^{\frac{2n}{d}}>T_{\text{train}}\text{,}$  for $\text{\quad for\ }n=d_{\text{extra}}/2,\cdots,d/2-1\text{.}$
@@ 3.3 Scaling Law for Larger Bases
Figure 7: The comparison between max practical context length and extrapolation upper bound predicted by Theorem 2. for RoPE tuned with larger bases on pre-training context. For bases larger than 800000, the extrapolation upper bound is beyond the evaluating context length.
@@ 3.4 Further Validation for Extrapolation
To further illustrate the causal relationship between critical dimension and extrapolation, we undertake the following three experiments. First, we set a max index of 4096 for the position embedding of the last 36 dimensions of $\bm{q}_{t},\bm{k}_{s}$ in LLaMA2 7B (Touvron et al., 2023b). We find that apart from an outlier in one attention head, the attention score variations are considerably reduced, leading to better extrapolation. Second, we visualize the attention score of the first 92 and the last 36 dimensions during extrapolation with Dynamic NTK (LocalLLaMA, 2023a) in Figure 7(a). Consistent with our theory, the attention scores of the last 36 dimensions display enhanced consistency compared with straightforward extrapolation. Lastly, we remove the final 36 dimensions of $\bm{q}_{t},\bm{k}_{s}$ in LLaMA2 7B (Touvron et al., 2023b) and fine-tuned using the setup in Appendix B.1. Remarkably, post-trimming fine-tuning substantially surpassed direct fine-tuning, allowing extrapolation beyond 20K tokens, as illustrated in Figure 7(b). This offers compelling evidence for the correlation between the critical dimension, attention score shifts, and extrapolation upper bound. It confirms that interpreting and enhancing RoPE-based LLM extrapolation from a periodic viewpoint is both valid and effective.
@@
In summary, we establish a comprehensive framework from a periodic perspective and identify the fundamental factors, the critical dimension for RoPE-based extrapolation. This not only unveils the root causes of attention score explosion highlighted in Chen et al. (2023) and Han et al. (2023), but also intuitively demonstrates how adjusting the base during fine-tuning can elevate the extrapolation capabilities of RoPE-based LLMs. Furthermore, we answer Q3 in the Introduction and introduce the Extended Scaling Law of RoPE-based Extrapolation in Appendix C, the combination of Theorem 1., Lemma 1. and Theorem 2., for tuning in a longer context. Additionally, we also present the instructive value of our theory for other works focused on achieving longer context in the testing phase in Appendix D. Finally, as depicted in Figure 1, tuning RoPE with a base of 500 or 1000000 can both outperform Linear PI (Chen et al., 2023) and NTK method (LocalLLaMA, 2023b; a). Besides, we compare the performance of LLaMA2 7B Touvron et al. (2023b) with base 500 and 1000000 in 1M context length in Table 1 and find that for extrapolating to unpredictable length, RoPE with base 500, has remarkable advantages, especially when combined with log-scaled attention (Su, 2023b).
@@ 4 Related Work
Recently, there have been a great deal of efforts devoted to expanding the context length of RoPE-based LLMs to 16K, 32K, or even 100K (Chen et al., 2023; LocalLLaMA, 2023b; Pal et al., 2023). Considering that LLMs, such as LLaMA2 (Touvron et al., 2023b), have already acquired sufficient knowledge in the pre-training stage and demonstrated excellence in short-context tasks, the emerging extrapolation improvements have primarily focused on the fine-tuning and the testing phase.
Extrapolating RoPE in Fine-tuning
In the fine-tuning phase, current strategies primarily concentrate on using smaller rotary angles to depict long context features. For example, Chen et al. (2023) extends the context window to 16K by linearly reducing the rotary angles with $\lambda=T_{\text{extra}}/T_{\text{train}}$ to align the input position index within the original context size. Rozière et al. (2023) the achieve 100K context window for code by increasing the rotary base to 1000000 and further training with 16K context length. Besides, Pal et al. (2023) also extends the context window to 32K by a new truncation strategy for refining the down-sampling method, which even sets some rotary angles to zero. Additionally, Peng et al. (2023) proposes a scaling function, fine-tunes LLaMA2 with 64K context, and finally achieves the 128K context length. Effective as they are, these methods all need additional tuning in longer contexts or face a disastrous collapse after the extrapolation bound.
Extrapolating RoPE in Inference
In the inference phase, context window division or restriction is a well-discussed strategy for any LLM without further training. Effective as it is, it prohibits tokens of their deserved global attention (Han et al., 2023) or destroys the order between token chunks (Ratner et al., 2022; Su, 2023a), which is unsupported for long document summarization (Su, 2023a). Besides, it is worth noting the remarkable effect of the Neural Tangents Kernel (NTK) method (LocalLLaMA, 2023b), especially the dynamic version (LocalLLaMA, 2023a) as follows:
for $\text{for }\bm{A}_{t,s}\text{, }\theta_{n}={\left(10000\cdot\alpha_{t}\right)}^{-2n/d}\text{, where }\alpha_{t}=\max\left(1,2^{\left\lceil\log_{2}{\frac{t}{T_{\text{train}}}}\right\rceil+1}-1\right)\text{.}$ (7)
Dynamic NTK  [^2] In this work, we follow the form of scaling function in QWenLM (Alibaba, 2023) simply decreases the rotary angle exponentially with scaling coefficient $\alpha$ as a function w.r.t. the inference length, which is discussed in detail in Appendix D. Dynamic NTK enables LLaMA2(Touvron et al., 2023b) to extrapolate without any further tuning or window assisting and keeps good performance within 16K context. However, Dynamic NTK still presents an obvious degradation in performance around 32K context size as shown in Figure 1.

### v2

@@ Abstract
The extrapolation capability of Large Language Models (LLMs) based on Rotary Position Embedding (Su et al., 2021) is currently a topic of considerable interest. The mainstream approach to addressing extrapolation with LLMs involves modifying RoPE by replacing 10000, the rotary base of $\theta_{n}={10000}^{-2n/d}$ in the original RoPE, with a larger value and providing longer fine-tuning text. In this work, we first observe that fine-tuning a RoPE-based LLM with either a smaller or larger base in pre-training context length could significantly enhance its extrapolation performance. After that, we propose Scaling Laws of RoPE-based Extrapolation, a unified framework from the periodic perspective, to describe the relationship between the extrapolation performance and base value as well as tuning context length. In this process, we also explain the origin of the RoPE-based extrapolation issue by critical dimension for extrapolation. Besides these observations and analyses, we achieve extrapolation up to 1 million context length within only 16K training length on LLaMA2 7B and 13B (Touvron et al., 2023b).
@@ 1 Introduction
Large Language Models (LLMs) have become the dominant architecture in a variety of natural language processing tasks(OpenAI, 2023; Touvron et al., 2023a; b), while Transformers (Vaswani et al., 2017) based on Rotary Position Embedding (RoPE) (Su et al., 2021) have become the dominant backbone in wide range of LLM design (Chowdhery et al., 2022; Nijkamp et al., 2023; Touvron et al., 2023a; b). While RoPE can theoretically represent sequences through trigonometric functions, as detailed in Appendix A, its performance drops when the input sequence or context length surpasses the training length(Press et al., 2022; Chen et al., 2023), seen in Figure 1. This extrapolation problem (Press et al., 2022) limits tasks like long text modeling and summarization(An et al., 2023).
@@
Figure 1: Perplexity on Books3 (Presser, 2020) with different extrapolation methods, including Dynamic NTK (bloc97, 2023b). RoPE fine-tuned with a smaller or larger base on the original training length of 4K or a much longer context of 16K, could outperform other extrapolation strategies and extrapolate to 100K context length.
Concerning the extrapolation issue with RoPE, different works have provided various interpretations and corresponding solving attempts. These works could divided into two schools of thought. One limits the scope of self-attention (Ratner et al., 2022; Han et al., 2023) given the fact that RoPE-based self-attention fails to keep stable beyond training context and exhibits attention score explosion as well as monotonous entropy increase (Chen et al., 2023; Han et al., 2023). The other aims to capture longer contexts with smaller rotation angles and longer fine-tuning context (Chen et al., 2023; Peng et al., 2023). Current popular methods, such as Dynamic NTK (bloc97, 2023a) and Code LLaMA (Rozière et al., 2023), mainly come from the second approach. Both approaches adapt RoPE to longer contexts with a larger rotary base. Specifically, Dynamic NTK (bloc97, 2023a) adjusts the base with a coefficient increasing with the length of inference, allowing RoPE-based LLMs to adapt simultaneously to longer context, while Code LLaMA (Rozière et al., 2023) directly sets the base at 1000000 and gets further trained on 16K context length, yielding a context beyond 100K.
@@
In this paper, we conduct further experiments on increasing and decreasing the rotary base in Section 2 and subsequently discover that adjusting the rotary base in both directions can contribute to the extrapolation of RoPE-based LLMs. Building upon these observations, we provide a comprehensive explanation for the seemingly counter-intuitive phenomenon from a periodic perspective. Meanwhile, we establish a unified theoretical framework for RoPE-based extrapolation known as the Scaling Laws of RoPE-based Extrapolation [^1] It is important to note that the scaling laws proposed in this work are irrelevant with the well-known scaling laws (Kaplan et al., 2020). In this paper, scale refers to the adjustment of RoPE’s rotary angles (bloc97, 2023a), rather than the change of the model size. Unlike the scaling laws (Kaplan et al., 2020) that are empirically derived, our scaling laws define mathematical relations between context window size and rotary base, supported by experiments.. We pinpoint specific changes during base reduction that lead to a significant boost in extrapolation in Section 3.1 and identify the upper bound of extrapolation for larger bases in Section 3.3. This clarifies how Code LLaMA (Rozière et al., 2023) manages a 100K extrapolation with only 16K training context. Furthermore, we validate our theories in Section 3.4 and Appendix C, shedding light on both the core principles of Dynamic NTK (bloc97, 2023a) and pinpointing instability sources self-attention computations in RoPE-based extrapolation (Han et al., 2023). Finally, we present the contributions and guiding significance of this work for other methods that achieve extrapolation during the inference phase. In summary, our contributions are as follows, and codes are available at https://github.com/OpenLMLab/scaling-rope.
@@ 2.1 Larger Bases Promise Better Extrapolation
(a) LLaMA2 7B
(b) LLaMA2 13B
(c) LLaMA2 7B
(d) LLaMA2 13B
First, Larger bases allow LLaMA2 (Touvron et al., 2023b) to extrapolate beyond its training context length, aligning with findings from Rozière et al. (2023). Secondly, the extrapolation with larger bases has a clear limit where language modeling perplexity stays consistent. Beyond this limit, the extrapolation performance declines significantly. Furthermore, as the base value rises, LLaMA2 can extrapolate to a longer context. Finally, compared to Dynamic NTK (bloc97, 2023a), RoPE tuned with larger bases degrades much quicker beyond its extrapolation upper bound. Therefore, for fine-tuning with larger bases, the performance beyond the upper bound could be consistently overtaken by Dynamic NTK. Nevertheless, within the upper bound, this approach still outperforms Dynamic NTK by a considerable margin, leading to a context beyond 100K with only a 4K tuning length, when the base is set over 600000.
2.2 Smaller Bases Also Promise Better Extrapolation
We then conduct the extrapolation experiments with smaller bases, using the same setup as for larger bases. Interestingly, even though this goes against common research findings (bloc97, 2023a; Rozière et al., 2023), fine-tuning RoPE with smaller bases on the original context length still boosts extrapolation, as shown in Figure 3. It also extends the context window beyond the training length. Yet, there are distinct differences when comparing RoPE with smaller bases to larger ones.
Firstly, RoPE with smaller bases does not have a distinct upper bound of extrapolation. While perplexity worsens as context length grows, this decline is gentler with smaller bases. Secondly, the enhancement in extrapolation of RoPE with smaller bases is not uniform. Between a base of 10000 and 8000, the extrapolation performance exhibits a tiny improvement. Then between a base of 8000 and 2608, improvement is moderate. After that, from 2608 to 1304 and further to 652, the improvement becomes more pronounced. Finally, when the rotary base is 500, the extrapolation curve becomes sufficiently smooth thus resulting in strong extrapolation over 48K context length and superior performance over Dynamic NTK (bloc97, 2023a).
@@ 2.2 Smaller Bases Also Promise Better Extrapolation
Figure 4: Perplexity and accuracy on the validation data from Books3 dataset (Presser, 2020) of LLaMA2 7B and 13B (Touvron et al., 2023b). In (a) and (b), the first row shows perplexity while the second row shows accuracy, both measured cumulatively. The first column shows the change of extrapolation performance w.r.t. rotary base at different context windows with a line plot, while the second column visualizes it with a heat map. In the heat map of perplexity, the value is log-scaled.
@@ 3.2 Critical Dimension for Extrapolation
$\bm{A}_{t,s}&=\mathrm{Re}\left[\underbrace{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\sum_{n=0}^{d/2-1}\tilde{q}_{t}^{(n)}\tilde{k}_{s}^{(n)}{}^{*}e^{i(t-s)\theta_{n}}}_{\text{full attention scores in RoPE}}\right]\\ &=\mathrm{Re}\left[{\underbrace{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{n=0}^{d_{\text{extra}}/2-1}\tilde{q}_{t}^{(n)}\tilde{k}_{s}^{(n)}{}^{*}e^{i(t-s)\theta_{n}}}_{\text{reliable part for extrapolation}}+\underbrace{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sum_{n=d_{\text{extra}}/2}^{d/2-1}\tilde{q}_{t}^{(n)}\tilde{k}_{s}^{(n)}{}^{*}e^{i(t-s)\theta_{n}}}_{\text{OOD part for extrapolation}}}\right]\end{aligned}\text{.}$ (2)
@@
$T_{n}=\frac{2\pi}{\theta_{n}}=2\pi\cdot{10000}^{\frac{2n}{d}}\leq T_{\text{train}}\text{,}$  $\text{\quad for\ }n=0,\cdots,d_{\text{extra}}/2-1\text{,}$ (3)
$T_{n}=\frac{2\pi}{\theta_{n}}=2\pi\cdot{10000}^{\frac{2n}{d}}>T_{\text{train}}\text{,}$  $\text{\quad for\ }n=d_{\text{extra}}/2,\cdots,d/2-1\text{.}$
@@ 3.3 Scaling Law for Larger Bases
(c) Validation on LLaMA2 70B
Figure 7: The comparison in 100K context length and 7B/13B model between max practical context length and extrapolation upper bound predicted by Theorem 2. for RoPE tuned with larger bases on pre-training context. We also validate our theory in 32K context length and 70B size and find that the tuning points of the perplexity curve match the theoretical upper bounds of length extrapolation.
@@ 3.4 Further Validation for Extrapolation
To further illustrate the causal relationship between critical dimension and extrapolation, we undertake the following three experiments. First, we set a max index of 4096 for the position embedding of the last 36 dimensions of $\bm{q}_{t},\bm{k}_{s}$ in LLaMA2 7B (Touvron et al., 2023b). We find that apart from an outlier in one attention head, the attention score variations are considerably reduced, leading to better extrapolation. Second, we visualize the attention score of the first 92 and the last 36 dimensions during extrapolation with Dynamic NTK (bloc97, 2023a) in Figure 8(a). Consistent with our theory, the attention scores of the last 36 dimensions display enhanced consistency compared with straightforward extrapolation. Lastly, we remove the final 36 dimensions of $\bm{q}_{t},\bm{k}_{s}$ in LLaMA2 7B (Touvron et al., 2023b) and fine-tuned using the setup in Appendix B.1. Remarkably, post-trimming fine-tuning substantially surpassed direct fine-tuning, allowing extrapolation beyond 20K tokens, as illustrated in Figure 8(b). This offers compelling evidence for the correlation between the critical dimension, attention score shifts, and extrapolation upper bound. It confirms that interpreting and enhancing RoPE-based LLM extrapolation from a periodic viewpoint is both valid and effective.
@@
In summary, we establish a comprehensive framework from a periodic perspective and identify the fundamental factors, the critical dimension for RoPE-based extrapolation. This not only unveils the root causes of attention score explosion highlighted in Chen et al. (2023) and Han et al. (2023), but also intuitively demonstrates how adjusting the base during fine-tuning can elevate the extrapolation capabilities of RoPE-based LLMs. Furthermore, we answer Q3 in the Introduction and introduce the Extended Scaling Law of RoPE-based Extrapolation in Appendix C, the combination of Theorem 1., Lemma 1. and Theorem 2., for tuning in a longer context. Additionally, we also present the instructive value of our theory for other works focused on achieving longer context in the testing phase in Appendix D. Finally, as depicted in Figure 1, tuning RoPE with a base of 500 or 1000000 can both outperform Linear PI (Chen et al., 2023) and NTK method (bloc97, 2023b; a). Besides, we compare the performance of LLaMA2 7B Touvron et al. (2023b) with base 500 and 1000000 in 1M context length in Table 1 and find that for extrapolating to unpredictable length, RoPE with base 500, has remarkable advantages, especially when combined with log-scaled attention (Su, 2023b).
@@ 4 Related Work
Long sequence modeling is an important issue in LLM applications. It faces the following three challenges: the quadratic computational complexity, the difficulty of parallelism for sequence length, and the collapse of the performance beyond the training length. First of all, due to the quadratic complexity of the self-attention mechanism in Transformers, long sequence modeling requires a large computational overhead. Early efficient Transformers reduced computational overhead through methods such as recalculation like ReFormer(Kitaev et al., 2020) and sparse attention such as LongFormerBeltagy et al. (2020) and BigBird(Zaheer et al., 2020). After that, LinFormer(Wang et al., 2020) linearizes the self-attention complexity through low-rank approximation, and Performer(Choromanski et al., 2021) expands the softmax calculation through a kernel function. Compared with these efficient attention work based on approximation, FlashAttention(Dao et al., 2022; Dao, 2023) reduces HBM reading and writing and turns self-attention into a form that can be iteratively calculated in a loop, which greatly improves the utilization of computing resources and has become the standard for long-sequence modeling.
The second challenge is parallelism for sequence length, that is, Sequence Parallelism (SP) (Li et al., 2023c). Since the memory and calculation of a single device are limited, ColossalAI-SP (Li et al., 2023c) first proposes the segmentation and parallelism for the sequence dimension in addition to tensor parallelism for the hidden dimension and pipeline parallelism for model depth. On this basis, Ring Attention(Liu et al., 2023) uses blockwise self-attention to split long sequences into different devices and overlap the communication of key-value blocks. Besides, LightSeq(Li et al., 2023b) further improves the efficiency of long sequence modeling through load balancing for causal language modelings and a re-materialization-aware checkpointing strategy. Although the above sequence parallelism can achieve infinitely long sequence modeling, they ignore the compatibility with the existing efficient self-attention mechanisms such as FlashAttention(Dao et al., 2022; Dao, 2023) to achieve input with almost infinity. On the contrary, Megatron-SP(Korthikanti et al., 2022) only uses sequence parallelism during Dropout and Layernorm operations, thereby reducing activation redundancy. In addition, DeepSpeed-Ulysses(Jacobs et al., 2023) uses an alternative all-to-all collective communication gathering weight for attention computation when segmenting the sequence, avoiding communication overhead that originally increases with length.
Recently, there have been a great deal of efforts devoted to solving the collapse of the performance beyond the training length and expanding the context length of RoPE-based LLMs to 16K, 32K, or even 100K (Chen et al., 2023; Su, 2023c; bloc97, 2023b). Considering that LLMs, such as LLaMA2 (Touvron et al., 2023b), have already acquired sufficient knowledge in the pre-training stage and demonstrated excellence in short-context tasks, the emerging extrapolation improvements have primarily focused on the fine-tuning and the testing phase. Initially, the context window is a well-discussed strategy for any LLM without further training. Effective as it is, it prohibits tokens of their deserved global attention (Han et al., 2023) or destroys the order between token chunks (Ratner et al., 2022; Su, 2023a), which is unsupported for long document summarization (Su, 2023a). Contrastively, Chen et al. (2023) extends the context window to 16K by linearly reducing the rotary angles with $\lambda=T_{\text{extra}}/T_{\text{train}}$ to align the input position index within the original context size. Besides, it is worth noting the remarkable effect of the Neural Tangents Kernel (NTK) method (bloc97, 2023b), especially the dynamic version (bloc97, 2023a)  [^2] In this work, we follow the form of scaling function in QWenLM (Alibaba, 2023) simply decreases the rotary angle exponentially with scaling coefficient $\alpha$ as a function w.r.t. the inference length, which is discussed in detail in Appendix D. Dynamic NTK enables LLaMA2(Touvron et al., 2023b) to extrapolate directly and keeps good performance within 16K context length. However, Dynamic NTK still presents an obvious degradation in performance around 32K context size as shown in Figure 1. Inspired by NTK methods, Rozière et al. (2023) achieves 100K context window for code by increasing the rotary base to 1000000 and further training with 16K context length. Besides, Pal et al. (2023) also extends the context window to 32K by a new truncation strategy for refining the down-sampling method, which even sets some rotary angles to zero. Additionally, Peng et al. (2023) proposes a scaling function, fine-tunes LLaMA2 with 64K context, and finally achieves the 128K context length. Effective as they are, these methods still face a disastrous collapse after an unpredictable extrapolation bound.